OpenStudy (anonymous):

-2)^3 - (6)(-5) - square root of 49 I keep getting -29 and apparently the answer is 15. What am I doing wrong?

Directrix (directrix):

(-2)^3 - (6)(-5) - square root of 49 = -8 + 30 - 7 = 30 - 15 = 15. -------- Please post your work. Thanks.

OpenStudy (anonymous):

Okay so -8 + 30 is 22 - 7 is 15. Ugh I don't know why I kept getting -22. I ALWAYS mix up my signs and I have a test tomorrow!!

Directrix (directrix):

@annaf1018 --> Practice on this problem - same instructions as above. 4 - (-2)^3 - (5)(-1)^3 = ? Test your signs here also: (-1)^2 -1^2 -(-1)^2